OpenStudy (anonymous):

greatest common factor of 14a^3b, 8a^2b^2 and 22ab^3

jhonyy9 (jhonyy9):

i think 2ab but not is sure

OpenStudy (anonymous):

factor each expression: 14a^3b=2*7*a*a*a*b 8a^2b^2=2*2*2*a*a*b*c 22ab^3=11*2*a*b*b*b find the factors that all 3 share they share 2ab.